OpenStudy (anonymous):

Is this statement true? "If U, W are subspaces of \(\mathbb{R}^n\) and dim U + dim W = n, then \(U\cap W = {0}\)"

OpenStudy (anonymous):

It should be false. :( Here is an counter-example. Suppose a basis of U is \(\{e_1\}\), and a basis of W is \(\{e_1, e_2, ...,e_{n-1}\}\), then dim U = 1, dim W = n-1, dim U +dim W = 1 + (n-1) = n. However, \(U\cap W = \{e_1\}\)

OpenStudy (anonymous):

\[\Large \checkmark \]

OpenStudy (anonymous):

Thanks :)